4. Prove that for positive integer n we have 169|33n+3-26n-27. 


1
Expert's answer
2022-06-28T03:20:27-0400

Let "P(n)" be the proposition that "169|3^{3n+3}-26n-27" for the positive integer "n."

Basis Step

"P(1)" is true because "3^{3(1)+3}-26(1)-27=676" is divisible by "169."

Inductive Step

Assume that "P(k)" is true for an arbitrary nonnegative integer "k." That is, we assume that "3^{3k+3}-26k-27" is divisible by "169."

We must show that when we assume that "P(k)" is true, then "P(k + 1)" is also true. That is, we must show that "3^{3(k+1)+3}-26(k+1)-27" is divisible by "169."

We find that


"3^{3(k+1)+3}-26(k+1)-27"

"=27(3^{3k+3})-26k-26-27"

"=27(3^{3k+3}-26k-27)+26(26k)+26(27)-26"

"=27(3^{3k+3}-26k-27)+676(k+1)"

"676" is divisible by "169,"and "3^{3k+3}-26k-27" is divisible by "169" by the inductive hypothesis. Then "27(3^{3k+3}-26k-27)+676(k+1)" is divisible by "169." This completes the inductive step of the proof.

We have completed the basis step and the inductive step. Hence, by mathematical induction "P(n)" is true for all positive integers "n."

That is, we have proved that "169|3^{3n+3}-26n-27" for the positive integer "n."